OpenStudy (anonymous):

find a solution to the initial value problem Y" +Y' = 0 y(pi/3) = 2, y'(pi/3) = -4

OpenStudy (anonymous):

\(y(x)=A\sin(x)+B\cos(x)\) \(y'(x)=A\cos(x)-B\sin(x)\) \(y(\pi/3)=A\sin(\pi/3)+B\cos(\pi/3)=2\) \(y'(\pi/3)=A\cos(\pi/3)iB\sin(\pi/3)=-4\) Solve the last 2 equations for A and B

OpenStudy (anonymous):

Sorry, the last equation should be: y′(π/3)=Acos(π/3)-Bsin(π/3)=−4

OpenStudy (anonymous):

I have to use the method concerning characteristic equation and the general solution, \[y = c1e^(t \alpha) \cos(\beta t) + c2e^(t \alpha)\sin(\beta t)... where... r = \alpha + or - \beta i\]

OpenStudy (anonymous):

nevermind, I made a copy error and the second y is not primed...already figured it out. Thanks for help tho!